OpenStudy (anonymous):

Help?? In economics, the Gini index of inequality (G), is one of the ways to measure the degree of concentration in the distribution of resources among members of a population. Is given by: \[G = 2\int\limits_{0}^{1}[x-L(x)]dx\] The L (x) function is called Lorentz curve and Gini index (G) always has a value between 0 and 1, so that the smaller the index, the more uniform the distribution of income, and the higher the index , will be the most concentrated income in the hands of a few individuals. If, for a given population, the Lorentz curve is L(x) = 0,8^2+0,2x, calculate the Gini ind

OpenStudy (kendricklamar2014):

I dont know, sorry :(

OpenStudy (anonymous):

thank you

OpenStudy (kendricklamar2014):

Sorry :(

OpenStudy (rock_mit182):

don´t ypu have to only integrate the function

OpenStudy (mathmale):

So, you want to integrate\[G = 2\int\limits\limits_{0}^{1}[x-L(x)]dx~?\]

OpenStudy (mathmale):

and is your L(x) equal to L(x) = 0,8^2+0,2x, ? \[ L(x) = 0.8^2+0.2x~?\]

OpenStudy (rock_mit182):

\[G =2\int\limits_{0}^{1} [x-(ax ^{2}+bx)]dx\]

OpenStudy (rock_mit182):

use the liner properties of an integral

OpenStudy (mathmale):

@rock_mit182 : Yes, and thank you for your contributions!! But please wait until the person who posted this question, fabio, responds, before you type in any more.

OpenStudy (rock_mit182):

ok

OpenStudy (rock_mit182):

:/

OpenStudy (mathmale):

@fabiomartins ?

OpenStudy (anonymous):

calculate the Gini index, do not know how to do

OpenStudy (mathmale):

Please see what @rock-mit182 has typed in so far. Let us know what part of his/her contributions you understand and which you don't.

OpenStudy (mathmale):

@fabiomartins ? We're waiting for your response.

OpenStudy (anonymous):

I did not understand, because I'm Brazilian, the result of this calculation is what you posted?

OpenStudy (mathmale):

Fabio! Your not understanding has little or nothing to do with your being Brazilian. this "Gini index" problem looks awful at first, but once you realize that you have L(x) and substitute that into the integral, the problem becomes a relatively easy one.

OpenStudy (mathmale):

Earlier, Rock_mit182 typed in:\[G =2\int\limits\limits_{0}^{1} [x-(ax ^{2}+bx)]dx\]and this is the correct thing to do. Just replace his (ax^2 + bx) with L(x) = 0.8^2+0.2x.

OpenStudy (anonymous):

It is possible to do step by step?

OpenStudy (mathmale):

Then\[G =2\int\limits\limits\limits_{0}^{1} [x-(ax ^{2}+bx)]dx\]becomes \[G =2\int\limits\limits\limits_{0}^{1} [x-(0.8x^2+0.2x)]dx\]

OpenStudy (mathmale):

Can you now integrate this?

OpenStudy (rock_mit182):

eu falo um poquinho portugues e voce ?

OpenStudy (mathmale):

I'm doing the best I can to simplify the problem for you. Combine 1.00x and 0.2x within the parentheses. What do you get for your integrand?

OpenStudy (anonymous):

sou Brasileiro, muito complicado esse cálculo, não estou entendendo

OpenStudy (mathmale):

Afraid not, Fabio. Not many people in the USA speak Portuguese. :( If you don't undrstand something I type in English, ask for clarification, and I will type it in different words.

OpenStudy (anonymous):

I'm trying

OpenStudy (mathmale):

Eu posso usar o Google Translate para escrever para você em Português. Tanto quanto possível, vamos nos focar na matemática e não na língua.

OpenStudy (rock_mit182):

ok lembra uma coisa, em matematicas voce tem que ter claro os conceptos

OpenStudy (rock_mit182):

depois tudo é proceso

OpenStudy (mathmale):

Vou deixar vocês dois trabalham juntos, porque parece que vocês dois podem se comunicar em Português. Deixe-me saber por mensagem privada, se você precisar da minha ajuda ainda mais.

OpenStudy (anonymous):

O cálculo de integral quando é só o cálculo eu até entendo, mas dessa forma que está ai, não consigo fazer..

OpenStudy (rock_mit182):

tudo legal ? @mathmale eu acho que noss precisamos de voce , i think we're going to need your help

OpenStudy (anonymous):

Com certeza precisamos sim do mathmale, muito gentil você, e atencioso...muito obrigado por vocês estarem me ajudando...

OpenStudy (rock_mit182):

tudo bem eu vou a falar um poquinho de integral é o basica pra voce enteder

OpenStudy (anonymous):

ok

OpenStudy (rock_mit182):

voce conhece a notação

OpenStudy (mathmale):

\[G =2\int\limits\limits\limits\limits_{0}^{1} [x-(0.8x^2+0.2x)]dx \]

OpenStudy (mathmale):

becomes\[G =2\int\limits\limits\limits\limits_{0}^{1} [1.2x-0.8x^2)]dx\]

OpenStudy (mathmale):

acordo?

OpenStudy (anonymous):

certo

OpenStudy (rock_mit182):

L é a integral da função , no intervalo entre 0 e 1 dois (2) é uma constante. entao voce ñao precisa de ela pra integar

OpenStudy (mathmale):

\[G =2\int\limits\limits\limits\limits\limits_{0}^{1} [1.2x-0.8x^2)]dx~\rightarrow~\]

OpenStudy (anonymous):

deixa eu ver se entendi....

OpenStudy (mathmale):

\[G =2\int\limits\limits\limits\limits\limits_{0}^{1} [1.2]dx2\int\limits\limits\limits\limits\limits_{0}^{1} (-0.8x^2)]dx\]

OpenStudy (mathmale):

acordo?

OpenStudy (anonymous):

ok, vou mandar ver se está correto...

OpenStudy (mathmale):

\[G =2\int\limits\limits\limits\limits\limits\limits_{0}^{1} [1.2]dx~+~2(-0.8)\int\limits\limits\limits\limits\limits\limits_{0}^{1} (x^2)]dx\]

OpenStudy (rock_mit182):

voce podi separa as dois funcoes, por a propiedade da operador linear

OpenStudy (mathmale):

Estou fazendo meus cálculos manualmente enquanto, parece que você está fazendo com uma calculadora online. Você precisa comparar seus resultados e os meus. Deixe-me saber se você encontrar discrepâncias.

OpenStudy (anonymous):

Pode continuar mathmale

OpenStudy (anonymous):

rock?

OpenStudy (mathmale):

Agora é hora de você meu amigo Fabio de "mostrar o seu material" .... mostre o que você mesmo pode fazer!

OpenStudy (anonymous):

Então, gostaria de saber se está correto o que postei?

OpenStudy (mathmale):

\[G =2.4\int\limits\limits\limits\limits\limits\limits\limits_{0}^{1} dx~-1.6\int\limits\limits\limits\limits\limits\limits\limits_{0}^{1} (x^2)]dx\]

OpenStudy (mathmale):

Embora eu não estou indo para verificar todos os detalhes de seu post para você, parece que você está definitivamente fazendo a coisa certa aqui. Parabéns.

OpenStudy (anonymous):

O meu resultado delu G=0,26

OpenStudy (mathmale):

Meu resultado é 1,8667. Você fez alguma subtração? Isso pode ser por isso que os resultados são diferentes. Eu não iria gastar muito mais tempo com este problema, especialmente porque você tem muitos outros para trabalhar.

OpenStudy (mathmale):

É maravilhoso que possamos nos comunicar usando o Google Translate. Tente você mesmo em algum momento: translate.google.com. Deixe-me saber se você precisar de ajuda com o Google Translate. Onde você mora no Brasil?

OpenStudy (anonymous):

Sim, e você?

OpenStudy (anonymous):

tem facebook?

OpenStudy (mathmale):

Eu vivo no sul da Califórnia, ao norte de Los Angeles. Estive em Brasil 3x.

OpenStudy (anonymous):

muito bom, não conheço outro país, você tem facebook?

OpenStudy (mathmale):

Eu não uso Facebook muito. Mas dê-me o seu nome no Facebook e eu posso enviar-lhe uma mensagem de que maneira.

OpenStudy (anonymous):

https://www.facebook.com/eng.fabiomartins